What is the empirical formula for a compound composed of 0.1417 mol of carbon (C), 0.0708 mol of hydrogen (H), and 0.0354 mol of nitrogen (N)?


1
Expert's answer
2022-02-27T22:53:03-0500

Moles= mass/rfm

Moles of carbon= 0.1417/12= 0.0118

Moles of hydrogen= 0.0708/1= 0.0708

Moles of nitrogen= 0.0354/14= 0.0025

Multiplying by a common factor ×1000

C= 11.8

H= 70.8

N= 2.5

Formula= C12H71N3